# AutoCodeRover: Autonomous Program Improvement

### v1

@@ 4.3. Stratified Context Search
The set of context retrieval APIs listed in Table 1 serves as building blocks for search relevant code context. With the context retrieval APIs and the LLM-idenfitied keyword “hints” from the problem statement, a set of possible API invocations can be derived by using the identified keywords as API parameters. We discuss a few observations on using these API invocations to gather code context, and propose a stratified context retrieval process.
Our first observation is that, if the context retrieval starts from a single API invocation, the LLM agent can get “stuck” into a “local optimum” context after making a sequence of API invocations. For example, in the issue django-13230 mentioned in Figure 4, if the retrieval starts from the invocation search_method("add_item"), the implementation of method add_item can be considered by the LLM agent as a sufficient context, since it appears to be relevant to the problem statement. However, in such cases, this context alone is not sufficient for the LLM to reason about the root cause. In this problem, the root cause is an error in the usage instead of definition of add_item. On the other hand, if all API invocations are executed at once, a large code context can be retrieved, especially when the problem statement mentions many class and method names. This large code context can be difficult for an LLM to comprehend, or may even exceed its context window.
@@
With these two observations, we propose a stratified search process for invoking context retrieval APIs, as illustrated in Figure 5. From a problem statement, stratified search iteratively invokes retrieval APIs to gather project code context, and finally outputs a list of potentially buggy locations to be fixed. In each stratum, we prompt the LLM agent to select a set of necessary API invocations, based on the current context. In stratum 1, the current context only contains the problem statement; in the following strata, the context contains both the problem statement and the code searched so far. By allowing LLM to select more than one API invocations and also instructing it to only select the necessary ones, we make best use of the context, building what we deem an optimal context.
After the API invocations in a stratum are executed, the newly retrieved code snippets are added to the current context. The LLM agent is then prompted to analyze whether the current context is sufficient for understanding the issue, thereby deciding whether (a) we continue the iterative search process to another stratum, or (b) we decide on the buggy locations which will then be considered for fixing.
@@ 4.4. Analysis-Augmented Context Retrieval
Apart from designing an autonomous system that retrieves code context from a Github issue description, we also investigate how program analysis and debugging techniques can augment the workflow. Specifically, we integrate the Spectrum-based Fault Localization (SBFL) analysis into AutoCodeRover to study the effect of such a test-based dynamic analysis in an autonomous program improvement framework. In this investigation, we make a test-suite $T$ available to AutoCodeRover, in addition to the problem statement $P$ and codebase $C$.
@@
The goal of SBFL is to identify the location of software faults (Wong et al., 2016b). Given a test-suite $T$ containing both passing and failing tests, SBFL considers control-flow differences in the passing and failing test executions, and assigns a suspiciousness score to different program locations. This suspiciousness score can be computed with various metrics such as Tarantula (Jones et al., 2002) and Ochiai (Abreu et al., 2007). Program elements (e.g. statements/basic blocks) with the highest suspiciousness scores are identified as likely fault locations. SBFL can be performed at different granularities of program elements, such as statements or basic blocks. Since the LLM can process reasonably long code snippets, we use method-level SBFL in AutoCodeRover. Given an test-suite, SBFL can be used to directly output a few program locations to be repaired. However, the accuracy of SBFL highly relies on the quality of the test-suite (Keller et al., 2017) - since the SBFL results are effectively an abstraction of the differential of control flows between passing tests and failing tests. Therefore, instead of replacing the stratified context retrieval with SBFL, we use the SBFL identified locations to augment the search process. Before AutoCodeRover enters the context retrieval stage, we provide the SBFL-identified locations to the LLM agent as “results from an external analysis tool that identifies suspicous code”. The main role of SBFL-identified locations is to reveal more “hints” on relevant classes and methods beyong those mentioned in the problem statement. The LLM agent can then use the context retrieval APIs to examine these program locations. Since the SBFL-identified locations are presented to the agent together with the problem statements, the agent can then cross-reference between these two sources of information. For example, if one of the SBFL-identified method names is more closly related to the problem statement, the LLM is more likely to invoke the search_method API on this name. We will demonstrate this observation in Section 6.2.
@@ 4.5. Patch Generation
Beyond performing patch generation solely from an LLM, there are also potentials in augmenting the patch generation stage with traditional techniques used in program repair. For example, we can use repair anti-patterns (Tan et al., 2016) to rule out patches that are syntactically correct but resemble a few patterns that are likely to be incorrect. We can also obtain a ranking of the patches in the patch pool based on various criteria such as (a) not satisfying anti-patterns (b) patch size so that smaller patches are preferred and, (c) optionally, satisfying the tests in the test-suite which were used for spectrum-based fault localization (SBFL). All of these are validation activities that are often carried out on a pool of patch candidates in program repair research (Goues et al., 2019).
@@ 5. Experiment Setup
We evaluate AutoCodeRover in a recently proposed benchmark SWE-bench lite (Jimenez et al., 2024), comprising 300 real-life GitHub issues from 11 popular repositories. The only input is the natural language description in the original GitHub issue and its corresponding buggy codebase. Details of SWE-bench lite appear in Section 2.2.
@@
We selected two recent LLM-based agent systems Swe-agent (Yang et al., 2024) and Magis (Tao et al., 2024) designed for SWE-bench tasks as baselines and compare their performance against our AutoCodeRover. Swe-agent is publicly available as a GitHub repository [^4] https://github.com/princeton-nlp/SWE-agent, so we replicated it as Swe-agent-rep with the default setting based on the provided scripts. In contrast to Swe-agent, we do not have access to Magis, so we take the most relevant reported result from their technical report (Tao et al., 2024). To avoid the natural randomness of LLM, we repeat our experiments with AutoCodeRover and Swe-agent-rep on SWE-bench lite three times. We use (1) the percentage of resolved instances, (2) average time cost, and (3) average token cost to evaluate the effectiveness of the tools. The three evaluation metrics represent overall effectiveness, time efficiency, and economic efficacy in resolving real-world GitHub issues. For AutoCodeRover and Swe-agent-rep, we report the average and total number of resolved instances across three repetitions.
@@
We use the state-of-the-art OpenAI GPT-4 (gpt-4-0125-preview) as the foundation inference model for both AutoCodeRover and Swe-agent-rep. In AutoCodeRover, the GPT-4 model is responsible for selecting search APIs to retrieve codebase context, refining the issue description, and writing a final patch. For parameters of GPT-4, we set a low temperature=0.2, max_tokens=1024 to produce relatively deterministic results and enable sufficient reasoning length for AutoCodeRover, and all other parameters remain default. For Swe-agent-rep, we increased the default cost limit of interacting with GPT-4 model when resolving an instance to two USD and kept other settings as default. Note that, both two tools do not have a time limit, AutoCodeRover terminates either a patch is generated or the context retrieval stage repeats ten times. Swe-agent-rep terminates either a patch is generated or the cost of resolving an issue reaches two USD. Experiment results and artifacts for AutoCodeRover and Swe-agent-rep can be found at https://github.com/nus-apr/auto-code-rover.
@@
All experiments are conducted on an x86_64 Linux server with Ubuntu 20.04 installed. Note that, in Table 3 and Figure 6, we only report the result of Swe-agent-rep on 274 task instances. This is because Swe-agent-rep fails to run on the remaining 26 instances (mostly matplotlib) in our environment, due to potential system-level dependency differences.
@@ 6.1. RQ1: Overall Effectiveness on SWE-bench
Table 2. Overall Result of AutoCodeRover (ACR) and baselines on SWE-bench lite. ”-” indicates data not available.
@@
Reported result on SWE-bench subset
Magis (Tao et al., 2024) 10.28% ( - ) - -
Devin (Labs, 2024) 13.86% ( - ) ¿ 600 -
Reported result on SWE-bench lite
Swe-agent (Jimenez et al., [n. d.]) 17.00% (51) 93 70181 (\$0.741) [^5] This cell denotes the average cost per instance in our replication experiment (SWE-bench lite)
ACR-1 16.00% (48) 159 36780 (\$0.426)
ACR-2 15.67% (47) 166 38664 (\$0.448)
ACR-3 16.67% (50) 195 37362 (\$0.430)
@@
ACR-all 22.33% (67) - -
Result on SWE-bench lite, assuming test-suite available
@@
We first measure the overall effectiveness of AutoCodeRover and baselines with the number of resolved task instances in SWE-bench lite. With the goal of understanding to what extent the current AI systems can automatically resolve real-life software issues, the only inputs we provided are the natural language issue description and a local code repository checked out at the erroneous version. We repeated the experiment of AutoCodeRover on SWE-bench lite three times, and the results are denoted as ACR-{1-3}. (For brevity, we use ACR to denote AutoCodeRover in the table.) In addition to that, we also report the average and total number of resolved software issues across the three runs, the results are denoted as ACR-avg and ACR-all respectively. The baseline results of Magis (Tao et al., 2024), Devin (Labs, 2024) are evaluated in another subset of SWE-bench, which we cannot directly compare with, so we take the results from their technical reports as a reference.
The result in Table 2 indicates that AutoCodeRover can resolve 15.67%-16.67% task instances in SWE-bench lite (300 tasks in total), which is on par with the reported results from the best baseline tool Swe-agent (Jimenez et al., [n. d.]). The number varies due to the randomness of large language models. On average, AutoCodeRover can resolve up to 16.11% task instances in SWE-bench lite without any human assistance. We also investigated the union of all resolved tasks in the three runs, in which the percentage of resolved task instances increased to 22.33%, showing the remarkable capability of AutoCodeRover and its potential to be deployed in the future.
Comparision with Swe-agent
We replicated Swe-agent with two USD as cost budget for conversation with LLM per task instance in our environment (Swe-agent-rep, terminates either a patch is generated or reaches the two USD budget), as the tool is publicly available. Specifically note that, out of the 300 tasks, there are 26 task instances that failed to be run in Swe-agent-rep. This might be due to the difference in the system environments, so we report the comparison of both tools on the remaining 274 task instances in Table 3 and Figure 6 separately. On average, AutoCodeRover resolved 17.15% task instances, whereas Swe-agent-rep resolved 10.83%. When considering the union of all resolved tasks across three repetitions, AutoCodeRover resolved 23.73% out of the 274 tasks, whereas Swe-agent-rep resolved 16.06%. We further analyzed the commonly and uniquely resolved instances between AutoCodeRover and Swe-agent-rep in Figure 6, and we found that AutoCodeRover and Swe-agent-rep complement each other in different scenarios. AutoCodeRover uniquely resolved 30 task instances, which benefited from the fine-grained code context search at the AST level to precisely locate the bug locations (e.g., django-13401 searches three necessary methods at one time). On the other hand, the main reason that AutoCodeRover failed on the 9 unique instances resolved by Swe-agent-rep is the confusion made by unimplemented search APIs (e.g., search_file invoked in django-12286). In such cases, AutoCodeRover can generate invalid search results when unimplemented APIs are invoked. This implies more robust search APIs are desired in future work to improve the results further.
Time / Token Cost
We are also interested in assessing the feasibility of deploying AutoCodeRover in the real world in terms of time and economic cost. On average, AutoCodeRover takes 173 seconds and 37602 tokens (equivalent to 0.435 USD) to resolve one task instance in SWE-bench lite. In comparison, our replication experiments with Swe-agent-rep cost 70181 tokens (equivalent to 0.741 USD) per task instance. Interestingly, it costs  2.77 days for developers to create pull requests for the 57 task instances that are also resolved by AutoCodeRover, and the other 10 task instances take even longer to be closed by developers (ranging from 34 - 4023 days). The short response time and low cost show the significant potential for AutoCodeRover to play as a first guard in future autonomous software maintenance.
Table 3. The percentage of resolved task instances of AutoCodeRover and Swe-agent-rep on the 274 instances of SWE-bench lite.
% of Resolved Task AutoCodeRover Swe-agent-rep
Repetition 1 17.15% (47) 10.22% (28)
Repetition 2 16.42% (45) 12.04% (33)
Repetition 3 17.88% (49) 10.22% (28)
Average 17.15% (47) 10.83% (29.7)
Total 23.72% (65) 16.06% (44)
Figure 6. The Venn diagram represents the intersection of all resolved task instances between AutoCodeRover and Swe-agent-rep on the 274 instances of SWE-bench lite.
6.2. RQ2: Effect of SBFL
The last two rows in Table 2 show the results of ACR-val and ACR-val-sbfl. With the additional information provided by SBFL, the number of resolved tasks increased from 51 to 61 (i.e. 17.00% to 20.33% resolved rate on SWE-bench lite). Compared to the validation-only setting, additionally adding the SBFL component helps to resolve 20 unique task instances. Moreover, when comparing with task instances resolved in all other runs combined (i.e. ACR-{1-3} and ACR-val), ACR-val-sbfl still uniquely resolves 7 task instances that are not resolved in any of the other runs.

### v2

@@ 4.3. Stratified Context Search
The set of context retrieval APIs listed in Table 1 serves as building blocks for searching relevant code context. With the context retrieval APIs and the LLM-identified keyword “hints” from the problem statement, a set of possible API invocations can be derived by using the identified keywords as API parameters. We discuss a few observations on using these API invocations to gather code context, and propose a stratified context retrieval process.
Our first observation is that the context retrieval should not be restricted to a single API invocation. For example, in the issue django-13230 mentioned in Figure 4, if the retrieval starts from the invocation search_method("add_item"), the implementation of method add_item can be considered by the LLM agent as a sufficient context, since it appears to be relevant to the problem statement. However, searching from only one method can lead to incomplete context for the agent to reason about root cause of the problem. On the other hand, if all API invocations are executed at once, a large code context can be retrieved, especially when the problem statement mentions many class and method names. This large code context can be difficult for an LLM to comprehend, or may even exceed its context window.
@@
With these two observations, we propose a stratified search process for invoking context retrieval APIs, as illustrated in Figure 5. From a problem statement, stratified search iteratively invokes retrieval APIs to gather project code context, and finally outputs a list of potentially buggy locations to be fixed. In each stratum, we prompt the LLM agent to select a set of necessary API invocations, based on the current context. In stratum 1, the current context only contains the problem statement; in the following strata, the context contains both the problem statement and the code searched so far. By allowing LLM to select more than one API invocations and also instructing it to only select the necessary ones, we make the best use of the context, building what we deem an optimal context.
After the API invocations in a stratum are executed, the newly retrieved code snippets are added to the current context. The LLM agent is then prompted to analyze whether the current context is sufficient for understanding the issue, thereby deciding whether (a) we continue the iterative search process, or (b) we decide on the buggy locations which will be considered for fixing.
@@ 4.4. Analysis-Augmented Context Retrieval
We also investigate how program debugging techniques can augment our workflow. Specifically, we integrate the Spectrum-based Fault Localization (SBFL) analysis into AutoCodeRover to study the effect of such a test-based dynamic analysis. We make a test-suite $T$ available to AutoCodeRover, in addition to the problem statement $P$ and codebase $C$.
@@
The goal of SBFL is to identify the location of software faults (Wong et al., 2016b). Given a test-suite $T$ containing both passing and failing tests, SBFL considers control-flow differences in the passing and failing test executions, and assigns a suspiciousness score to different program locations. This suspiciousness score can be computed with various metrics such as Tarantula (Jones et al., 2002) and Ochiai (Abreu et al., 2007). Program elements (e.g. statements/basic blocks) with the highest suspiciousness scores are identified as likely fault locations. SBFL can be performed at different granularities of program elements, such as statements or basic blocks. Since the LLM can process reasonably long code snippets, we use method-level SBFL in AutoCodeRover. Given an test-suite, SBFL can be used to directly output a few program locations to be repaired. However, the accuracy of SBFL highly relies on the quality of the test-suite (Keller et al., 2017) - since the SBFL results are effectively an abstraction of the differential of control flows between passing tests and failing tests. Therefore, instead of replacing the stratified context retrieval with SBFL, we use the SBFL identified methods to augment the search process. Before AutoCodeRover enters the context retrieval stage, we provide the SBFL-identified methods to the LLM agent as “results from an external analysis tool that identifies suspicous code”. The main role of SBFL-identified methods is to reveal more “hints” on relevant classes and methods beyond those mentioned in the problem statement. The LLM agent can then use the context retrieval APIs to examine these methods. Since the SBFL-identified methods are presented to the agent together with the problem statements, the agent can then cross-reference between these two sources of information. For example, if one of the SBFL-identified method names is more closly related to the problem statement, the LLM is more likely to invoke the search_method API on this name. We will demonstrate this observation in Section 6.2.
@@ 5. Experiment Setup
We evaluate AutoCodeRover in recently proposed benchmarks SWE-bench and SWE-bench lite (Jimenez et al., 2024b), comprising 2294 and 300 real-life GitHub issues, respectively. The only input is the natural language description in the original GitHub issue and its corresponding buggy codebase. Details of SWE-bench and SWE-bench lite appear in Section 2.2.
@@
We selected two LLM-based agent systems Devin (Labs, 2024) and Swe-agent (Yang et al., 2024) as baselines and compare their performance against AutoCodeRover. Swe-agent is publicly available as a GitHub repository [^3] https://github.com/princeton-nlp/SWE-agent, so we replicated it as Swe-agent-rep in SWE-bench lite with the default setting based on the provided scripts. In contrast to Swe-agent, we do not have access to Devin, so we take the most relevant reported result from their technical report (Team, 2024). To avoid the natural randomness of LLM, we repeat our experiments with AutoCodeRover and Swe-agent-rep three times when possible, and report the average and total numbers across three repetitions. We use (1) the percentage of resolved instances, (2) average time cost, and (3) average token cost to evaluate the effectiveness of the tools. These evaluation metrics represent overall effectiveness, time efficiency, and economic efficacy in resolving real-world GitHub issues.
@@
We use the state-of-the-art OpenAI GPT-4 (gpt-4-0125-preview) as foundation inference model for both AutoCodeRover and Swe-agent-rep. In AutoCodeRover, the GPT-4 model is responsible for selecting search APIs to retrieve codebase context, refining the issue description, and writing a final patch. For parameters of GPT-4, we set a low temperature=0.2, max_tokens=1024 to produce relatively deterministic results and enable sufficient reasoning length for AutoCodeRover, and all other parameters remain as per default. For Swe-agent-rep, we increased the default cost limit of interacting with GPT-4 model (for resolving a task instace) to two USD and kept other settings as default. Note that, both two tools do not have a time limit, AutoCodeRover terminates either a patch is generated or the context retrieval stage repeats ten times. Swe-agent-rep terminates either a patch is generated or the cost of resolving an issue reaches two USD. Experiment results and artifacts for AutoCodeRover and Swe-agent-rep can be found at https://github.com/nus-apr/auto-code-rover.
@@
All experiments are conducted on an x86_64 Linux server with Ubuntu 20.04 installed.
@@ 6.1. RQ1: Overall Effectiveness on SWE-bench
We first measure the overall effectiveness of AutoCodeRover and baselines with the number of resolved task instances in SWE-bench. With the goal of understanding to what extent the current AI systems can automatically resolve real-life software issues, the only inputs we provided are the natural language issue description and a local code repository checked out at the erroneous version. We repeated the experiment of AutoCodeRover three times and presented the average and total number of resolved software issues across the three runs. The average and total results are denoted as ACR-avg and ACR-all respectively (for brevity, we use ACR to denote AutoCodeRover in this section). When reporting time and token/cost for ACR-all, we report the time and cost required for running each task three times. Since Devin was evaluated on a random 25% subset of SWE-bench (Labs, 2024), we also report results of AutoCodeRover on this subset (we refer to this as “SWE-bench Devin subset”). Table 2 shows the overall result in full SWE-bench, SWE-bench Devin subset, SWE-bench lite respectively. Figure 6 shows a visual summary of AutoCodeRover’s comparison with Swe-agent and Devin.
Results on SWE-bench.
In the full SWE-bench, ACR-all (union of from the three ACR runs) resolved 15.95% task instances taking a time of 701 seconds per task (less than 12 minutes). In comparison, Swe-agent resolved 12.29% tasks in full SWE-bench, according to their report (Jimenez et al., 2024a) (we did not replicate Swe-agent-rep on the full SWE-bench due to the high cost).
Comparison with Devin.
The state-of-the-art closed-source baseline tool Devin (Labs, 2024) is evaluated in a random 25% subset of SWE-bench. To compare AutoCodeRover with Devin, we report AutoCodeRover’s result on the 570 task instances Devin was evaluated on. The results of AutoCodeRover are taken from the ACR runs on full SWE-bench. In the SWE-bench Devin subset, the union of three runs of AutoCodeRover successfully resolved 15.79% of the task instances, which is higher than Devin. Figure 7(a) provides a more detailed exposition of the resolved tasks. Besides, the times taken by AutoCodeRover and Devin are comparable.
Results on SWE-bench lite.
We performed another round of experiments with AutoCodeRover on SWE-bench lite (300 instances). The results reported in Table 2 indicate that on average ACR-avg can resolve 16.11% task instances in SWE-bench lite, which is at par with the reported results from Swe-agent (Jimenez et al., 2024a). We also investigated the union of all resolved tasks in the three runs, in which the percentage of resolved task instances increased to 22.33%.
Table 2. Overall Result of AutoCodeRover (ACR) and baselines on full SWE-bench, SWE-bench Devin subset, and SWE-bench lite. ”-” indicates data not available.
@@
Reported result on full SWE-bench (size=2294)
Swe-agent (Jimenez et al., 2024a) 12.29% (282) 93 -
ACR-avg 10.59% (243) 234 40177 (\$0.464)
ACR-all 15.95% (366) 701 120530 (\$1.392)
Reported result on SWE-bench Devin subset (size=570)
Devin (Labs, 2024) 13.86% (79) ¿ 600 -
ACR-avg 10.47% (59.7) 231 39068 (\$0.452)
ACR-all 15.79% (90) 692 117204 (\$1.357)
Reported result on SWE-bench lite (size=300)
Swe-agent (Jimenez et al., 2024a) 17.00% (51) 93 69976 (\$0.739) [^4] This cell denotes the average cost per instance in our replication experiment (SWE-bench lite)
@@
ACR-all 22.33% (67) 520 112806 (\$1.304)
Figure 6. Summary of Results
Detailed Comparison with Swe-agent
Since Swe-agent is publicly available, we also attempted to run Swe-agent on SWE-bench lite. We replicated Swe-agent with two USD as cost budget for conversation with LLM per task instance in our environment (denote as Swe-agent-rep, it terminates either a patch is generated or reaches the two USD budget). Table 3 shows that when considering the union of all resolved tasks across three repetitions, AutoCodeRover resolved 22.33% out of the 300 tasks, whereas Swe-agent-rep resolved 14.67%. We further analyzed the commonly and uniquely resolved instances between AutoCodeRover and Swe-agent-rep in Figure 7(b), and we found that AutoCodeRover and Swe-agent-rep complement each other in different scenarios. AutoCodeRover uniquely resolved 32 task instances, which benefited from the fine-grained code context search at the AST level to precisely locate the bug locations (e.g., django-13401 searches three necessary methods at one time). On the other hand, the main reason that AutoCodeRover failed on the 9 unique instances resolved by Swe-agent-rep, is unimplemented search APIs (e.g., search_file invoked in django-12286). In such cases, AutoCodeRover can generate invalid search results when unimplemented APIs are invoked. This implies more robust search APIs are desired for future improvement. Since the evaluation environment is important to the result, to make the experimental replication complete and thorough, we also evaluated the generated patches of 300 task instances of SWE-bench lite within the Swe-agent docker environment which has been publicly released. However, Table 3 shows that the number of resolved issues for both ACR and Swe-agent-rep decreased, indicating that the docker released by SWE-bench team may need more work.
Time / Token Cost
We are also interested in assessing the feasibility of deploying AutoCodeRover in the real world in terms of time and economic cost. On average, AutoCodeRover takes 234 seconds and 40177 tokens (equivalent to 0.464 USD) to resolve one task instance in SWE-bench. In comparison, our replication experiments with Swe-agent-rep cost 69976 tokens (equivalent to 0.739 USD) per task instance. When considering the combined three repetitions, AutoCodeRover takes 701 seconds (~11.7 minutes) per task, which is below the 30-60 minute time limit deemed acceptable by developers for automated repair tools (Noller et al., 2022). Looking into the 67 issues resolved by AutoCodeRover in SWE-bench lite, it costs on average ~2.77 days for developers to create pull requests for 57 issues, and the other 10 issues take even longer to be closed by developers (ranging from 34 - 4023 days). The short response time and low cost show the significant potential for AutoCodeRover to act as a first step in future autonomous bug fixing.
Table 3. The percentage of resolved task instances of AutoCodeRover and Swe-agent-rep on SWE-bench lite. The column “In Swe-agent Docker” represents the evaluation result under the docker environment released publicly by Swe-agent team on 6th April 2024.
Replication result on SWE-bench lite (size=300)
In our environment In Swe-agent Docker
ACR Agent-Rep ACR Agent-Rep
Run 1 16.00% (48) 9.33% (28) 10.00% (30) 6.67% (20)
Run 2 15.67% (47) 11.00% (33) 10.33% (31) 7.00% (21)
Run 3 16.67% (50) 9.33% (28) 10.67% (32) 6.00% (18)
All 22.33% (67) 14.67% (44) 14.00% (42) 9.00% (27)
(a) AutoCodeRover and Devin on SWE-bench Devin subset.
(b) AutoCodeRover and Swe-agent-rep on SWE-bench lite.
Figure 7. Venn diagrams of resolved tasks instances.
Plausible / Correct patches.
Overfitting is a well-known challenge in the Automated Program Repair community (Gao et al., 2019). A program patch that passes the given test suite is said to be plausible. However, a plausible patch is deemed as overfitting if it fails to conform to the developer’s intent. Otherwise, it is deemed as correct. To further understand the patch quality of AutoCodeRover and baselines, we manually verify the correctness of task-resolving (i.e. plausible) patches in SWE-bench lite. Since three repetitions are performed, we consider a task to have a correct patch if any of the three repetitions produced a correct patch. A plausible patch is correct if it is semantically equivalent to the developer patch. In this verification process, at least two authors of the paper cross-validated each patch, and any disagreement was resolved with another author. Overall, on SWE-bench lite, ACR has a correctness rate of 65.7% (44 correct/67 plausible). Swe-agent-rep has a slightly higher correctness rate of 72.7% (32/44), but the absolute number of correctly resolved tasks is smaller than ACR. Finally, the correctness rate of Devin on SWE-bench Devin subset is 53.2% (42/79). We observed that vast majority of ACR’s overfitting patches (all but 2 of the overfitting patches) modify the same methods as the developer patches, but the code modifications are wrong. This means that even the overfitting patches from ACR are useful to the developer, since it helps in localization. The main causes of wrong modifications are the limits of the LLM’s capability or insufficient context. Apart from these, we noticed two other interesting causes of overfitting. One cause is that the issue creator gives a preliminary patch in the description. This patch can be different from the final developer patch, misleading the LLM. The other interesting cause is that the issue creator mentioned a case that needs to be handled. The LLM only fixes this mentioned case, but the developer fixed other similar cases as well. The two causes indicate that the issue description, just like test suites, can be an incomplete specification.
6.2. RQ2: Effect of SBFL
Table 4. Result of ACR-val, ACR-val-sbfl on SWE-bench lite, in one run only.
Tools Resolved Tasks Avg Time Avg Tokens
ACR-avg 16.11% (48.3) 173 37602 (\$0.435)
@@ 6.2. RQ2: Effect of SBFL
Results.
Table 4 shows that, with the additional information provided by SBFL, the number of resolved tasks increased from 51 to 61 (i.e. 17.00% to 20.33% resolved rate on SWE-bench lite). Compared to the validation-only setting, additionally adding the SBFL component helps to resolve 20 unique task instances. Moreover, when comparing with task instances resolved in all other runs combined (i.e. ACR-all and ACR-val) on SWE-bench lite, ACR-val-sbfl still uniquely resolves 7 task instances that are not resolved in any of the other runs.